OpenStudy (anonymous):

OK. I need help solving a linear system by using the elimination method, but I also need to arrange the like terms, but sadly I don't know how. The linear system is 1. 2x - y = 32 2. y -5x = 13. The second equation needs to be arranged. Please help?

OpenStudy (anonymous):

I'll try to help. Are you still around?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay, gimme a sec. (writing problems down)

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

The first trick, is to line the equations up, with all of the variables and numbers in columns. This makes it easier to see what we're looking at.

OpenStudy (anonymous):

Yeah I did that. I need to write it in standard form, but the x and y are in the different places.

OpenStudy (anonymous):

So, in other words, here is how I rewrote them. -y + 1.2x = 32.2 y - 5x = 13

OpenStudy (anonymous):

Oh, lol, I didn't even think about standard form.

OpenStudy (anonymous):

The original equation for the first one is 2x - y = 32

OpenStudy (anonymous):

How did you get 1.2x?

OpenStudy (anonymous):

Ahhhh! Okay. Let's rewrite the equations in standard form. ( I read your first post wrong)

OpenStudy (anonymous):

2x - y = 32 5x - y = 13

OpenStudy (anonymous):

Do you remember the additive identity property?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Okay. That property lets us move the addition/subtraction terms around however we like.

OpenStudy (anonymous):

That's how I was able to rearrange the terms

OpenStudy (anonymous):

Do you also remember that you can multiply both sides by a negative?

OpenStudy (anonymous):

I know that part to.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So did you arrange the like terms correctly?

OpenStudy (anonymous):

2x - y = 32 5x - y = 13 ?

OpenStudy (anonymous):

I'm going to multiply the top equation by negative one, and then I'm going to add the two equations together. And, yes.

OpenStudy (anonymous):

That's what I see.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Standard form is X + Y = Z

OpenStudy (anonymous):

Your going to solve for one variable right?

OpenStudy (anonymous):

Yup. One way we can accomplish this is by adding the two equations together.

OpenStudy (anonymous):

I thought standad form was Ax + By = c?

OpenStudy (anonymous):

Uhh.. Yes. You're more right. (Sorry, haven't used standard form in a while)

OpenStudy (anonymous):

It's cool.

OpenStudy (anonymous):

Okay. So let's multiply by (-1) 2x - y = 32 becomes -2x + y = -32

OpenStudy (anonymous):

Okay and then your going to add it?

OpenStudy (anonymous):

I got 3x = - 19

OpenStudy (anonymous):

Yes! Now, it becomes very apparent why adding equations is awesome. -2x + y = - 32 + 5x - y = 13 ( I need to double check the negatives)

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Yeah, it's -2x + y = -32 + 5x - y = -13 I didn't need to multiply the top equation by negative one, (oops!), so I multiplied the bottom one, so that they'll add up nicely again. (Sorry for the confusion)

OpenStudy (anonymous):

Would you add them together please?

OpenStudy (anonymous):

3x = - 45

OpenStudy (anonymous):

x = -15

OpenStudy (anonymous):

Yup! Now you just need to clean up.

OpenStudy (anonymous):

Now, substitute x for - 15 on another equation, and you'll get y. ^^

OpenStudy (anonymous):

-2(-15) + y = -32 30 + y = -32 y = -2

OpenStudy (anonymous):

Hmm..... Something went wrong. Argh! Sorry!!

OpenStudy (anonymous):

I use it in the original equation don't I?

OpenStudy (anonymous):

Alright, let's try this one more time! Hang on, gimme a second.

OpenStudy (anonymous):

2(-15) - y = 32 -30 - y = 32 -y = 62 y = -62

OpenStudy (anonymous):

Hey I checked..and your right.

OpenStudy (anonymous):

The solution is (-15, -62)

OpenStudy (anonymous):

O_O Yeah. I checked it. Just some bad 'rithmatic

OpenStudy (anonymous):

Ow. Okay, cool. Whew. I was worried I led you down the wrong path. Hehe

OpenStudy (anonymous):

It's cool.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Anyway, awesome. Take it easy, and God bless with the rest of your homework. ... Yikes. I think I need some sleep!

OpenStudy (anonymous):

Lol. Me too.